A store offers a mixture of raisins and dried pinapple. Raisins cost $3 per pound and dried pinapple cost $4 per pound. How many pounds are needed to make 20 pounds of mixture worth $3.60 per pound?
Answer by askmemath(368)

Let number of pounds of Raisins = R
Number of pounds of Pineapple = P
Total weight of mixture = 20
P+R = 20
price per pound of raisins = 3
Total price of raisins = 3 X R = 3R
Price per pound of pineapple = 4
Total price of Pineapple = 4 X P = 4P
We are making 20pounds of mixture worth 3.60 per pound
SO total price of the mixture = 20 x 3.6 = 72
So 3R + 4P = 72
R + P = 20
Can you solve from here?